OpenStudy (anonymous):

I'm stuck on this last question and i'm not sure what i'm doing wrong, I can't seem to find the right solution. 2x^2 - 3x - 4 = 0 From the Quadratic Formula -b ± √(b² - 4ac))/2a -3 ± √(-3^2 - 4(2)(4))/2(2) -3 ± √(-9 - 32)/4 (everything else above this point I think I did correctly, but I got stuck here, can someone help me?) -3 ± √(41)/4 -3 ± 6.4/4 This is what the solution should be: x ≈ 2.35 or x ≈ -0.85

OpenStudy (anonymous):

\[ x=\frac{1}{4} \left(3-\sqrt{41}\right)\\ y=\frac{1}{4} \left(3+\sqrt{41}\right) \]

OpenStudy (anonymous):

I'm afraid I still don't understand??

OpenStudy (anonymous):

@Renee99 b=-3 b^2 = 9 and not -9 This might be where you sent wrong.

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

Also -4( a c) =(-4) (2)(-4)= 32 and -32

OpenStudy (anonymous):

ohh, okay I see now, let me redo it and see if I get it right this time :)

OpenStudy (anonymous):

also -b =3 and not -3

OpenStudy (anonymous):

Use quadratic formula x = [-b + √(b² - 4ac)] / 2a or x = [-b - √(b² - 4ac)] / 2a So using the values a = 2 b = -3 c = -4 x = [-(-3) + √((-3)² - (4)(2)(-4)] / 2(2) or x = [-(-3) - √((-3)² - (4)(2)(-4)] / 2(2) so x = -[√(41)-3] / 4, or x = [√(41)+3] / 4

OpenStudy (anonymous):

Best answer me(:

OpenStudy (anonymous):

@Math4Life you should not ask for Best answer me. One should leav it to the asker.

OpenStudy (anonymous):

This is what I have now: -b ± √(b^2 - 4ac))/2a 3 ± √(-3^2 - 4(2)(-4))/2(2) 3 ± √(9 + 32)/4 3 ± √(41)/4 3 ± 10.25/4

OpenStudy (anonymous):

Yes, this is right.

OpenStudy (anonymous):

okay, so now how do I get to this: x ≈ 2.35 or x ≈ -0.85

OpenStudy (anonymous):

Use your calculator to compute \[x=\frac{1}{4} \left(3-\sqrt{41}\right)=-0.85\\ y=\frac{1}{4} \left(3+\sqrt{41}\right)=2.35 \]

OpenStudy (anonymous):

@eliassaab ok sorry

OpenStudy (anonymous):

but this is your whole problem if you want it @Renee99 Use quadratic formula x = [-b + √(b² - 4ac)] / 2a or x = [-b - √(b² - 4ac)] / 2a So using the values a = 2 b = -3 c = -4 x = [-(-3) + √((-3)² - (4)(2)(-4)] / 2(2) or x = [-(-3) - √((-3)² - (4)(2)(-4)] / 2(2) so x = -[√(41)-3] / 4, or x = [√(41)+3] / 4

OpenStudy (anonymous):

okay, phew.... gotta take a breather after that one, lol! Anyway, thanks for both yours help! I really appreciate it! :)